OpenStudy (anonymous):

Evaluate the integral by interpreting it in terms of areas. S^0 on the bottom -6 (5+root36-x^2)dx I have no idea how to work this problem out, please help!

ganeshie8 (ganeshie8):

Take a good look at the integrand, does it look familiar to you ?

OpenStudy (anonymous):

the -6 is on the bottom of the integral, um no not really

ganeshie8 (ganeshie8):

like this ?\[\large \int\limits_{-6}^0 \color{blue}{5+\sqrt{36-x^2}}\, dx\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

May be rearrange the integrand a bit \[\large y=\color{blue}{5+\sqrt{36-x^2}}\] \[\large (y\color{blue}{-5})^2=(\color{blue}{\sqrt{36-x^2}})^2\] \[\large \color{blue}{x^2}+(y\color{blue}{-5})^2=\color{blue}{36}\] what about now ? seen this before ?

OpenStudy (anonymous):

it has to do with circles right?

ganeshie8 (ganeshie8):